# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v3

@@ IV.2 Sub-Class B
$\mathcal{M}_{B}=\left(\begin{array}[]{cc}c_{1}Y^{a\over{q+1}}&c_{2}Y^{a\over{q+1}}+c_{3}Y^{b\over{q+1}}\\ c_{4}Y^{a\over{q+1}}+c_{5}Y^{b\over{q+1}}&c_{6}Y^{b\over{q+1}}\end{array}\right),~{}~{}$ (50)
@@
Conditions (52) may seem restrictive, but they actually constitute an underdetermined quadratic polynomial system (six variables, two quadratic equations, and one quadratic inequality constraint), and such systems are known to be either inconsistent (i.e., have no solution) or possess an infinite number of real and complex solutions [56]. In the present case, one can easily find a few families of solutions using simple trial and error, and hence the system is not inconsistent (i.e., there are an infinite number of solutions). It is worth noting that the effect of the quadratic inequality is only to impose some restrictions on the (infinity of) solutions to the two quadratic equations. The full set of solutions to this quadratic system can be determined (with the aid of computer algebra software) and expressed in parametric form but it is not necessary since the only mathematical requirement (for sub-class B to be well defined) is that the polynomial system not be inconsistent (i.e., the conditions on the constants can actually be satisfied).
IV.2.1 Illustrative Examples
@@
$\mathcal{M}^{(1)}_{B(I)}=\left(\begin{array}[]{cc}Y^{1\over{q+1}}&2Y^{1\over{q+1}}-{1\over 2}Y^{u(q)}\\ Y^{1\over{q+1}}+{1\over 2}Y^{u(q)}&Y^{u(q)}\end{array}\right),~{}~{}~{}$ (53c)
@@
$\mathcal{M}^{(1)}_{B(III)}=\left(\begin{array}[]{cc}2Y^{v(q)}&Y^{v(q)}+2Y^{1\over{q+1}}\\ -Y^{v(q)}+5Y^{1\over{q+1}}&3Y^{1\over{q+1}}\end{array}\right),~{}~{}~{}$ (53f)
@@ IV.2.3 Generalizations of Cases I and III
$\mathcal{M}^{(\text{g})}_{B(I)}=\left(\begin{array}[]{cc}c_{1}Y^{1\over{q+1}}&c_{2}Y^{1\over{q+1}}+c_{3}V\\ c_{4}Y^{1\over{q+1}}+c_{5}V&c_{6}V\end{array}\right),~{}~{}~{}$ (61c)
@@
$\mathcal{M}^{(\text{g})}_{B(III)}=\left(\begin{array}[]{cc}c_{1}V&c_{2}V+c_{3}Y^{1\over{q+1}}\\ c_{4}V+c_{5}Y^{1\over{q+1}}&c_{6}Y^{1\over{q+1}}\end{array}\right).~{}~{}~{}$ (62c)
@@ IV.3.2 Significance of Sub-Class C
V Complex Matrices and Eigenvalues
@@
Model (6) admits complex matrices having real or complex eigenvalues. As long as condition (8) is satisfied, the Brody distribution holds for complex matrices and eigenvalues as well. An exhaustive treatment of all six possible complex cases will not be given here. For simplicity, only the two cases where the trace of $\mathcal{M}_{n}$ is real will be discussed. For purposes of illustration, it suffices to consider the complex generalization of sub-classes A and B and to give simple examples chosen from those sub-classes.
V.1 Complex Matrices with Real Eigenvalues
@@
Condition (31d) for sub-class A and the conditions in (52a)-(52c) for sub-class B each constitute an underdetermined quadratic polynomial system. As previously discussed, these systems possess an infinite number of real and complex solutions. In defining sub-classes A, B, and C, the constants $\{c_{j}\}$ were stipulated to be real. This was done to simplify the conditioning of the constants, but is not a strict requirement. If the constants $\{c_{j}\}$ are permitted to be complex, then the defining conditions can be generalized so as to admit complex matrices with real eigenvalues. For example, condition (31d) for sub-class A would be replaced by the following three conditions:
@@
$\mathcal{M}^{\text{nH}(1)}_{A}=\left(\begin{array}[]{cc}X+(3-i)Y^{1\over{q+1}}&\left(2+{1\over 2}i\right)Y^{-1}\\ \left(2+{1\over 2}i\right)Y^{q+3\over{q+1}}&X+(1+i)Y^{1\over{q+1}}\end{array}\right),~{}~{}$ (82)
@@
$\mathcal{M}^{\text{cs(1)}}_{B}=\left(\begin{array}[]{cc}Y^{a\over{q+1}}&{i\over 2}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)\\ {i\over 2}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)&-Y^{b\over{q+1}}\end{array}\right),~{}~{}~{}$ (85)
@@
V.2 Complex-Conjugate Eigenvalues

### v4

@@ V.2 Sub-Class B
$\mathcal{M}_{B}=\left(\begin{array}[]{cc}c_{1}Y^{a\over{q+1}}&c_{2}Y^{a\over{q+1}}+c_{3}Y^{b\over{q+1}}\\ c_{4}Y^{a\over{q+1}}+c_{5}Y^{b\over{q+1}}&c_{6}Y^{b\over{q+1}}\end{array}\right),$ (50)
@@
Conditions (52) may seem restrictive, but they actually constitute an underdetermined quadratic polynomial system (six variables, two quadratic equations, and one quadratic inequality constraint), and such systems are known to be either inconsistent (i.e., have no solution) or possess an infinite number of real and complex solutions [56]. In the present case, one can easily find a few families of solutions using simple trial and error, and hence the system is not inconsistent (i.e., there are an infinite number of solutions). It is worth noting that the effect of the quadratic inequality is only to impose some restrictions on the (infinity of) solutions to the two quadratic equations. The full set of solutions to this quadratic system can be determined (with the aid of computer algebra software) and expressed in parametric form but it is not necessary since the only mathematical requirement (for sub-class B to be well defined) is that the polynomial system not be inconsistent (i.e., the conditions on the constants can actually be satisfied). Some illustrative examples are given below.
V.2.1 Illustrative Examples
@@
$\mathcal{M}^{(1)}_{B(I)}=\left(\begin{array}[]{cc}Y^{1\over{q+1}}&2Y^{1\over{q+1}}-{1\over 2}Y^{u(q)}\\ Y^{1\over{q+1}}+{1\over 2}Y^{u(q)}&Y^{u(q)}\end{array}\right),$ (53c)
@@
$\mathcal{M}^{(1)}_{B(III)}=\left(\begin{array}[]{cc}2Y^{v(q)}&Y^{v(q)}+2Y^{1\over{q+1}}\\ -Y^{v(q)}+5Y^{1\over{q+1}}&3Y^{1\over{q+1}}\end{array}\right),$ (53f)
@@ V.2.3 Generalizations of Cases I and III
$\mathcal{M}^{(\text{g})}_{B(I)}=\left(\begin{array}[]{cc}c_{1}Y^{1\over{q+1}}&c_{2}Y^{1\over{q+1}}+c_{3}V\\ c_{4}Y^{1\over{q+1}}+c_{5}V&c_{6}V\end{array}\right),$ (61c)
@@
$\mathcal{M}^{(\text{g})}_{B(III)}=\left(\begin{array}[]{cc}c_{1}V&c_{2}V+c_{3}Y^{1\over{q+1}}\\ c_{4}V+c_{5}Y^{1\over{q+1}}&c_{6}Y^{1\over{q+1}}\end{array}\right).$ (62c)
@@ V.3.2 Significance of Sub-Class C
VI Simple Symmetric Models
Model (6) can also generate symmetric sub-classes of models. The extra condition of symmetry (in addition to condition (8)) is actually non-trivial since its imposition introduces algebraic constraints that often result in sets of conditions that have no fully real solutions. It can also happen that satisfaction of condition (8) precludes symmetry outright (e.g., sub-class B). Imposing symmetry reduces the number of distinct summand coefficients (i.e., the $\{k_{m}\}$ in model (6) or the constants $\{c_{j}\}$ in the context of sub-classes) and the reduction in algebraic degrees-of-freedom makes (non-trivial) satisfaction of condition (8) more difficult. Satisfaction of condition (8) can however be conveniently enforced by making use of a variable prefactor $\mathcal{A}$ that multiplies each of the matrix elements. (Recall that $\mathcal{A}=1$ for all sub-classes of models considered so far.) Any models that result from using such prefactors will (by definition) not be cropped, pruned, nor trimmed.
VI.1 Simple Low-Order Models: Examples
Among the simplest (non-trivial) models are the following symmetric sub-classes of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (respectively):
${1\over\sqrt{c^{2}(q)Y^{-{2\over{q+1}}}+1}}\left(\begin{array}[]{cc}c(q)&~{1\over 2}\left(c(q)+Y^{1\over{q+1}}\right)\\[4.26773pt] {1\over 2}\left(c(q)+Y^{1\over{q+1}}\right)&Y^{1\over{q+1}}\end{array}\right),$ (79c)
${1\over\sqrt{c^{2}(q)Y^{-{2\over{q+1}}}+1}}\left(\begin{array}[]{cc}c(q)&~\sqrt{c(q)\over{2}}Y^{1\over{2(q+1)}}\\[4.26773pt] \sqrt{c(q)\over{2}}Y^{1\over{2(q+1)}}&Y^{1\over{q+1}}\end{array}\right),$ (79f)
where $c(q)$ is a real (possibly $q$-dependent) constant that can be arbitrarily prescribed.
Models (79) can be immediately generalized as follows:
${1\over\sqrt{Y^{{2(a-1)\over{q+1}}}+Y^{{2(b-1)\over{q+1}}}}}\left(\begin{array}[]{cc}Y^{a\over{q+1}}&~{1\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)\\[4.26773pt] {1\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)&Y^{b\over{q+1}}\end{array}\right),~a\neq b$ (80c)
${1\over\sqrt{Y^{{2(a-1)\over{q+1}}}+Y^{{2(b-1)\over{q+1}}}}}\left(\begin{array}[]{cc}Y^{a\over{q+1}}&~{1\over\sqrt{2}}Y^{a+b\over{2(q+1)}}\\[4.26773pt] {1\over\sqrt{2}}Y^{a+b\over{2(q+1)}}&Y^{b\over{q+1}}\end{array}\right),~a\neq b$ (80f)
where the strictly unequal (and independent) power constants $\{a,b\}$ are real (possibly $q$-dependent) constants that can be arbitrarily prescribed (e.g., $\{a=\pi,b=\ln(q+{1\over 2})\}$). Models (80c) and (80f) are generally sub-classes of $\mathcal{M}_{2}$ and $\mathcal{M}_{3}$, respectively (the exceptions occur when one of $\{a,b\}$ is zero in which case (80c) and (80f) are sub-classes of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively).
A related model with zero trace is:
${1\over\sqrt{Y^{{a-2\over{q+1}}}+Y^{{b-2\over{q+1}}}}}\left(\begin{array}[]{cc}Y^{a\over{2(q+1)}}&Y^{b\over{2(q+1)}}\\ Y^{b\over{2(q+1)}}&~-Y^{a\over{2(q+1)}}\end{array}\right),~a\neq b$ (83)
which is a symmetric sub-class of $\mathcal{M}_{2}$ (unless one of $\{a,b\}$ is zero in which case (83) is a sub-class of $\mathcal{M}_{1}$).
VI.2 Model Construction Using Preset Polynomial Functions of $Y^{1\over{q+1}}$
There is no unique approach to constructing symmetric sub-classes of model (6). One could employ general algebraic approaches when considering the matrix-element summations (as was done in Section V), but one could also make use of specific predefined finite-sums with prescribed constants. To illustrate, consider (for simplicity) the case where the summand coefficients $\{k_{m}\}$ in model (6) are prescribed real numbers and the power constants $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ are positive integers; this corresponds to considering different polynomial functions of the fundamental random variable $Y^{1\over{q+1}}$. One could determine the general integer sets of values of $\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}\right\}$ that satisfy condition (8), which aligns with the general algebraic approach used in Section V, or one could alternatively prescribe the values of these constants (i.e., use specific preset polynomials) and then enforce the satisfaction of condition (8) through use of the matrix prefactor $\mathcal{A}$. The latter approach will be employed in the following two subsections.
The choice of polynomial(s) in the above-described construction is arbitrary. Polynomials obtained from binomial expansions are simple and well-suited due to their inherent symmetry. Sets of orthogonal polynomials are also convenient due to their many special properties and recurrence relations. Some specific examples are given in the following two subsections.
VI.3 Polynomial Constructions 1: Binomial Expansions
VI.3.1 Model 1
An example utilizing binomial expansions in the diagonal entries is the following general symmetric sub-class:
$\mathcal{A}\left(\begin{array}[]{cc}\left(c(q)+Y^{1\over{q+1}}\right)^{n}&~~\mathcal{P}\left(Y^{1\over{q+1}};c(q)\right)\\[4.26773pt] \mathcal{P}\left(Y^{1\over{q+1}};c(q)\right)&~\left(c(q)-Y^{1\over{q+1}}\right)^{n}\end{array}\right),$ (86)
where $n$ is a positive integer, $\mathcal{A}$ is a variable prefactor, and $\mathcal{P}$ is a polynomial in the variable $Y^{1\over{q+1}}$ (with $c(q)$-dependent coefficients). $\left\{\mathcal{A},\mathcal{P}\right\}$ are non-unique (for each fixed value of $n$) and can be freely prescribed subject to condition (8).
Example 1: one sub-class when $n=3$ is:
${1\over{\sqrt{9c^{4}+Y^{4\over{q+1}}}}}\left(\begin{array}[]{cc}\left(c+Y^{1\over{q+1}}\right)^{3}&~~3c^{2}Y^{1\over{q+1}}-Y^{3\over{q+1}}\\[4.26773pt] 3c^{2}Y^{1\over{q+1}}-Y^{3\over{q+1}}&~\left(c-Y^{1\over{q+1}}\right)^{3}\end{array}\right).$ (87c)
Example 2: one sub-class when $n=4$ is:
${1\over{c\sqrt{c^{4}+Y^{4\over{q+1}}}}}\left(\begin{array}[]{cc}\left(c+Y^{1\over{q+1}}\right)^{4}&~~4c^{3}Y^{1\over{q+1}}-4cY^{3\over{q+1}}\\[4.26773pt] 4c^{3}Y^{1\over{q+1}}-4cY^{3\over{q+1}}&~\left(c-Y^{1\over{q+1}}\right)^{4}\end{array}\right).$ (87f)
Note that the degree of $\mathcal{P}$ in (86) need not be equal to $n$. For instance, in Example 2 above (where $n=4$), $\mathcal{P}$ is a polynomial of degree $3$ (in the variable $Y^{1\over{q+1}}$).
VI.3.2 Model 2
An example utilizing binomial expansions in only one of the diagonal entries are the symmetric sub-classes:
$\mathcal{A}_{\pm}\left(\begin{array}[]{cc}\left(c(q)\pm Y^{1\over{q+1}}\right)^{n}&~~\mathcal{P}_{\pm}\left(Y^{1\over{q+1}};c(q)\right)\\[4.26773pt] \mathcal{P}_{\pm}\left(Y^{1\over{q+1}};c(q)\right)&~c(q)^{n}\pm Y^{n\over{q+1}}\end{array}\right),$ (90)
where (as in Model 1) $n$ is a positive integer, $\mathcal{A}_{\pm}$ are variable prefactors, and $\mathcal{P}_{\pm}$ are polynomials in the variable $Y^{1\over{q+1}}$. In each case (i.e., in the “++” or “-” case), $\left\{\mathcal{A}_{\pm},\mathcal{P}_{\pm}\right\}$ are non-unique (for each fixed value of $n$) and can be freely prescribed subject to condition (8).
Example 1: one “++” sub-class when $n=3$ is:
${1\over{c\sqrt{c^{2}+Y^{2\over{q+1}}}}}\left(\begin{array}[]{cc}\left(c+Y^{1\over{q+1}}\right)^{3}&~~{3\over 2}c^{2}Y^{1\over{q+1}}-{3\over 2}cY^{2\over{q+1}}\\[4.26773pt] {3\over 2}c^{2}Y^{1\over{q+1}}-{3\over 2}cY^{2\over{q+1}}&~c^{3}+Y^{3\over{q+1}}\end{array}\right).$ (91c)
Example 2: another “++” sub-class when $n=3$ is:
${1\over{c\left(c+Y^{1\over{q+1}}\right)}}\left(\begin{array}[]{cc}\left(c+Y^{1\over{q+1}}\right)^{3}&~~{3\over 2}c^{2}Y^{1\over{q+1}}+{3\over 2}cY^{2\over{q+1}}\\[4.26773pt] {3\over 2}c^{2}Y^{1\over{q+1}}+{3\over 2}cY^{2\over{q+1}}&~c^{3}+Y^{3\over{q+1}}\end{array}\right).$ (91f)
VI.3.3 Model 3
An example having zero trace is the general symmetric sub-class (of $\mathcal{M}_{n}$):
$\mathcal{A}_{n}\left(\begin{array}[]{cc}\left(c(q)-Y^{1\over{q+1}}\right)^{n}&~~\left(c(q)+Y^{1\over{q+1}}\right)^{n}\\[4.26773pt] \left(c(q)+Y^{1\over{q+1}}\right)^{n}&~-\left(c(q)-Y^{1\over{q+1}}\right)^{n}\end{array}\right),$ (94)
where $n$ is a positive integer.
Example: For $n=\{1,2,3\}$, the matrix prefactor $\mathcal{A}_{n}$ is:
$\mathcal{A}_{1}={1\over{2\sqrt{c^{2}Y^{-{2\over{q+1}}}+1}}},$ (95a)
$\mathcal{A}_{2}={1\over{2\sqrt{c^{4}Y^{-{2\over{q+1}}}+6c^{2}+Y^{2\over{q+1}}}}},$ (95b)
$\mathcal{A}_{3}={1\over{2\sqrt{c^{6}Y^{-{2\over{q+1}}}+15c^{4}+15c^{2}Y^{2\over{q+1}}+Y^{4\over{q+1}}}}}.$ (95c)
VI.3.4 Mixed Constructions
In constructing models using binomial expansions, it is not necessary for all matrix elements to be strictly comprised of only polynomial functions of $Y^{1\over{q+1}}$. The following are two simple low-order examples.
Example 1:
${1\over\sqrt{c^{2}+Y^{{2\over{q+1}}}}}\left(\begin{array}[]{cc}\left(c-Y^{1\over{q+1}}\right)^{2}&~\sqrt{2c}Y^{3\over{2(q+1)}}\\[4.26773pt] \sqrt{2c}Y^{3\over{2(q+1)}}&c^{2}-Y^{2\over{q+1}}\end{array}\right),$ (96c)
which is a sub-class of $\mathcal{M}_{3}$.
Example 2:
${1\over{c}\sqrt{c^{2}+Y^{{2\over{q+1}}}}}\left(\begin{array}[]{cc}\left(c-Y^{1\over{q+1}}\right)^{3}&~{3\over\sqrt{2}}\left(cY^{1\over{q+1}}\right)^{3/2}\\[4.26773pt] {3\over\sqrt{2}}\left(cY^{1\over{q+1}}\right)^{3/2}&c^{3}-Y^{3\over{q+1}}\end{array}\right),$ (96f)
which is a sub-class of $\mathcal{M}_{4}$.
Note that (ignoring matrix prefactors) the off-diagonal elements of models (96c) and (96f) are (apart from a constant) comprised of the same non-polynomial function of $Y^{1\over{q+1}}$ while their diagonal elements are comprised of (different) polynomial functions of $Y^{1\over{q+1}}$.
VI.3.5 Non-Polynomial Generalization
The binomial expansions $\left(c(q)\pm Y^{1\over{q+1}}\right)^{n}$ were used in models (86), (90), (94), and (96). One could also use the more general binomial expansions $\left(Y^{a\over{q+1}}\pm Y^{b\over{q+1}}\right)^{n}$, where $a\neq b$ are arbitrary real parameters. This would yield non-polynomial generalizations of the preceding constructions.
Example: Defining $C(a,b)\equiv\sqrt{Y^{2a\over{q+1}}+Y^{{2b\over{q+1}}}}$, the resulting generalizations of models (96c) and (96f) would be (respectively):
${1\over{C(a,b)Y^{b-1\over{q+1}}}}\left(\begin{array}[]{cc}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)^{2}&~\sqrt{2}Y^{a+3b\over{2(q+1)}}\\[4.26773pt] \sqrt{2}Y^{a+3b\over{2(q+1)}}&~Y^{2a\over{q+1}}-Y^{2b\over{q+1}}\end{array}\right),$ (97c)
which is a sub-class of $\mathcal{M}_{4}$, and
${1\over{C(a,b)Y^{a+b-1\over{q+1}}}}\left(\begin{array}[]{cc}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)^{3}&~{3\over\sqrt{2}}Y^{3(a+b)\over{2(q+1)}}\\[4.26773pt] {3\over\sqrt{2}}Y^{3(a+b)\over{2(q+1)}}&~Y^{3a\over{q+1}}-Y^{3b\over{q+1}}\end{array}\right),$ (97f)
which is a sub-class of $\mathcal{M}_{5}$.
VI.4 Polynomial Constructions 2: Orthogonal Polynomials
As mentioned in Section VI.2, model construction using predefined polynomials could also employ sets of orthogonal polynomials, including (but not limited to) Legendre, Hermite, Laguerre, and Chebyshev polynomials. Examples making use of the latter two polynomial sets are given below. In the following, the random variable $W\equiv Y^{1\over{q+1}}$ [c.f., Eq. (12)].
VI.4.1 Model 1: Laguerre Polynomials
Let $\mathcal{L}_{n}$ denote the Laguerre polynomial of degree $n$, where $n=\{0,1,2,3,\ldots\}$. The following model (among others) is obtained using the fundamental recurrence relation for Laguerre polynomials:
${1\over{\sqrt{1+(2n+1)^{2}W^{-2}}\mathcal{L}_{n}(W)}}\left(\begin{array}[]{cc}-\mathcal{L}_{n+1}(W)&~~{1\over 2}(W+2n+1)\mathcal{L}_{n}(W)\\[1.42271pt] {1\over 2}(W+2n+1)\mathcal{L}_{n}(W)&~n^{2}\mathcal{L}_{n-1}(W)\end{array}\right).$ (100)
Example: The case $n=2$ yields the model
$\mathcal{A}_{\mathcal{L}_{2}}(W)\left(\begin{array}[]{cc}W^{3}-9W^{2}+18W-6&~~{1\over 2}W^{3}+{1\over 2}W^{2}-9W+5\\[1.42271pt] {1\over 2}W^{3}+{1\over 2}W^{2}-9W+5&~-4W+4\end{array}\right),$ (101c)
where
$\mathcal{A}_{\mathcal{L}_{2}}(W)={1\over{\sqrt{1+25W^{-2}}\left(W^{2}-4W+2\right)}}.$ (101d)
VI.4.2 Model 2: Chebyshev Polynomials
Let $\{\mathcal{T}_{n},\mathcal{U}_{n}\}$ denote the Chebyshev polynomials of the first and second kinds (respectively) of degree $n$. The recurrence relations for each of these two polynomial sets and the various interrelationships between them allow the construction of many different symmetric models. For instance, one model is:
${1\over{\left(W+W^{-1}\right)\mathcal{U}_{n-1}(W)}}\left(\begin{array}[]{cc}W\mathcal{T}_{n}(W)&~~W\mathcal{U}_{n-1}(W)\\[0.85355pt] W\mathcal{U}_{n-1}(W)&~\mathcal{T}_{n+1}(W)\end{array}\right).$ (104)
Example: The case $n=2$ yields the model
${1\over{2\left(W^{2}+1\right)}}\left(\begin{array}[]{cc}2W^{3}-W&~2W^{2}\\[0.85355pt] 2W^{2}&~~4W^{3}-3W\end{array}\right).$ (107)
VII Complex Matrices and Eigenvalues
@@
Model (6) admits complex matrices having real or complex eigenvalues. As long as condition (8) is satisfied, the Brody distribution holds for complex matrices and eigenvalues as well. An exhaustive treatment of all six possible complex cases will not be given here. For simplicity, only the two cases where the trace of $\mathcal{M}_{n}$ is real will be discussed. For purposes of illustration, it suffices to consider the complex generalization of (the non-symmetric) sub-classes A and B and to give simple examples chosen from those sub-classes.
VII.1 Complex Matrices with Real Eigenvalues
@@
Condition (31d) for sub-class A, conditions (52a)-(52c) for sub-class B, and conditions (66a)-(66b) for sub-class C, each constitute an underdetermined quadratic polynomial system. As previously discussed, these systems possess an infinite number of real and complex solutions. In defining sub-classes A, B, and C, the constants $\{c_{j}\}$ were stipulated to be real. This was done to simplify the conditioning of the constants, but is not a strict requirement. If the constants $\{c_{j}\}$ are permitted to be complex, then the defining conditions can be generalized so as to admit complex matrices with real eigenvalues. For example, condition (31d) for sub-class A would be replaced by the following three conditions:
@@
$\mathcal{M}^{\text{nH}(1)}_{A}=\left(\begin{array}[]{cc}X+(3-i)Y^{1\over{q+1}}&\left(2+{1\over 2}i\right)Y^{-1}\\ \left(2+{1\over 2}i\right)Y^{q+3\over{q+1}}&X+(1+i)Y^{1\over{q+1}}\end{array}\right),$ (111)
@@
$\mathcal{M}^{\text{cs(1)}}_{B}=\left(\begin{array}[]{cc}Y^{a\over{q+1}}&{i\over 2}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)\\ {i\over 2}\left(Y^{a\over{q+1}}-Y^{b\over{q+1}}\right)&-Y^{b\over{q+1}}\end{array}\right),$ (114)
@@
VII.2 Complex-Conjugate Eigenvalues